A rectangle is 4 feet longer than it is wide. It's perimeter is 60 feet. Find the length an width of the rectangle

Answer by checkley79(3341)

L=W+4
P=2L+2W
60=2(W+4)+2W
60=2W+8+2W
60-8=4W
52=4W
W=52/4
W=13 ANS. FOR THE WIDTH.
L=13+4=17 ANS. FOR THE LENGTH.
PROOF:
60=2*17+2*13
60=34+26
60=60